Assuming two objects or two massages attracted to each other with a gravitational force of 16units. If the mass of both objects were tripped and the distance between the object doubled what will be the resulting force of attraction between the two objects


1
Expert's answer
2021-08-02T08:40:45-0400
"F'=\frac{3^2}{2^2}F\\F'=\frac{3^2}{2^2}16=36\ N"